OpenStudy (lexy148):

A sugar cookie recipe that makes 3 dozen cookies requires 2 1/2cups of powered sugar. if you want to make enough cookies so that each of 30 students in you child's class gets exactly two cookies, how much powdered sugar will you need? (can someone help me to put it into a proportion

OpenStudy (dan815):

well well well

OpenStudy (dan815):

this comes down to being able to cook really

OpenStudy (dan815):

3*12 cookies = 2.5 sugar

OpenStudy (dan815):

2 for each 30 kids mean u ned to make 60 cookies so 60 = (3*12)*60/((3*12)) = 2.5*(60/(3*12))

OpenStudy (dan815):

60 cookies = (3*12)*60/((3*12)) cookies = 2.5*(60/(3*12)) sugar

OpenStudy (dan815):

basically you see how many times more 60 is than 3 dozen and thats manytimes more sugar u need

OpenStudy (dan815):

3 dozen = 3*12 = 36 now.... 60 is 60/36 times more than 36 since 36 * 60/36=60 so we are going to need 60/36 times more sugar we need 2.5 sugar cups for 3dozen or 36 cookies so we need 2.5*60/36 sugar cups for 60 cookies